how many milliliters of a 10% hydrogen peroxide solution and how many milliliters of a 2% hydrogen peroxide solution should be mixed to obtain 500 mL if a 4% hydrogen peroxide solution?

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

how many milliliters of a 10% hydrogen peroxide solution and how many milliliters of a 2% hydrogen peroxide solution should be mixed to obtain 500 mL if a 4% hydrogen peroxide solution?
----
Equation:
hydro + hydro = hydro
0.02x + 0.10(500-x) = 0.04*500
------------------
2x + 10*500 - 10x = 4*500
----
-8x = -6*500
----
x = (3/4)500
x = 3*125 = 375 ml (amt. of 2% solution needed)
500-x = 125 ml (amt. of 10% solution needed)
--------------
Cheers,
Stan H.
--------------

Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6